OpenStudy (anonymous):

What is the exact diameter of an area rug with a circumference of 21 feet?

OpenStudy (diyadiya):

\[Circumference= 2 \times \pi \times Radius \]\[21= 2 \pi r\]Now find Radius Diameter is 2 x Radius

OpenStudy (anonymous):

is the answer 21 pi feet or 21/ pi feet

OpenStudy (diyadiya):

21/pi

OpenStudy (anonymous):

kk thx! :)

OpenStudy (diyadiya):

Welcome :)